How we can solve supply and demand equation for finding elasticity of demand and supply at equilibrium position
1
Expert's answer
2020-05-03T16:59:22-0400

We can write elasticity formula like this:

"E=\frac{\Delta Q}{\Delta P}\times\frac{P}{Q}."

Let's write demand formula:

Qd= aP+c.

Thus "\frac{\Delta Q}{\Delta P}= a" , bacease "a" shows the rise in quantity demanded if the price changes by 1.

Let (Q', P') be equilibrium point. Thus we can calculate elasticity of demand like:

"Ed=a\times \frac{P'}{Q'}" .

The same is with supply. If supply curve is Qs=bP+d, than the suplly elasticity will be:

"Es=b\times\frac{P'}{Q'}."